The table below shows the average scores from three classrooms after taking a self-report survey for test anxiety. The fall scores are pre-intervention, while the spring scores are post-intervention. Utilizing a paired sample t-test, decide if there is a significant difference between the fall and spring scores.

Perceived Anxiety Levels Around Literacy: Experimental Group

Classrooms

Fall

Spring

Classroom 4

23.8

21.6

Classroom 5

28.2

29.7

Classroom 6

26.2

30.3
Answer #1

From the given data,values of d = Fall - Spring are got as follows:

2.2, - 1.5, - 4.1

From d values, the following statistics are calculated:

n = 3

= - 1.1333

sd = 3.1660

SE = sd/

= 3.1660/ = 1.8279

Test statistic is given by:

t = - 1.1333/1.8279 = - 0.6200

ndf = 3 - 1 = 2

Take = 0.05

From Table, critical values of t = 4.3027

Since calculated value of t = - 0.6200 is greater than critical value of t = - 4.3027, the difference is not significant. Fail to reject null hypothesis.

Conclusion:

The data do not support the claim that there is a significant difference between the fall and spring scores.